If we solve X 2 =1 then we put X=+or- 1 Then if we solve x 8 =1, Will there also be plus or minus? Or only plus? Why?

If we solve
X2 =1 then we put 
X=+or- 1
Then if we solve x8=1,
Will there also be plus or minus?
Or only plus?
Why?

Grade:11

2 Answers

Aditya Gupta
5 years ago
If x belongs to real, then yes you are right in saying that there will be BOTH plus as well as minus. Because we can write x^8=1^8 as well as x^8= (-1)^8. Hence both are correct
Arun
5 years ago
Dear Jeet
 
in x8 = 1
 
x will be ± 1.
 
but if there is odd power then only one sign is used.
 
Regards
Arun